Circular No. 7212

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999cx and 1999cy J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by M. Wischnjewsky of two apparent supernovae on 20-min unfiltered T-Max films taken by L. E. Gonzalez with the University of Chile’s Maksutov telescope at Cerro El Roble:

SN  1999 UT  α_{2000}  δ_{2000}  Mag.  Offset
1999cx May 14.15 12^{h}07^{m}49.^{s}17 −20^{o}00′22′′.8 18.5 4′′ W, 7′′ S
1999cy May 13.36 11 18 49.47 + 0 37 08.7 18.0 7′′ W, 7′′ N

Confirmation of both supernovae was made by Hamuy on CCD B and V images obtained on May 19.14 and 19.13 UT with the European Southern Observatory 3.5-m New Technology Telescope. SN 1999cx is not visible on a film obtained with the Cerro El Roble telescope on Apr. 21.27 (limiting mag > 20), and SN 1999cy is not visible on a film taken on Apr. 16.09 (limiting mag > 20).

COMET C/1999 M2 (SOHO) D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports observations of another apparent Kreutz sungrazing comet that he discovered in images taken with the C2 coronagraph aboard SOHO. A short tail was visible. Astrometric measurements made by Biesecker (and reduced by G. V. Williams), together with orbital elements by B. G. Marsden, are given on MPEC 1999-M40.

1999 UT  α_{2000}  δ_{2000}
June 27.314  6^{h}17.^{m}3  +21^{o}50′

V364 PEGASI J. Y. Wei, J. Y. Hu, Y. L. Qiu, and Q. Y. Qiao, Beijing Astronomical

Observatory (BAO), report that V364 Peg (cf. IAUC 7159) was at unfiltered CCD magnitude 16.4 on June 11.77 UT.

NOVA IN NGC 6822 B. A. Skiff, Lowell Observatory, has posted a list of reliable photometric

comparison stars (12 < V < 21) within 5′ of this apparent nova (cf. IAUC

7211) at ftp://ftp.lowell.edu/pub/bas/starcats/n6822.seq.

1999 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green